Direct Materials Variances

Bellingham Company produces a product that requires 14 standard pounds per unit. The standard price is $11.5 per pound. If 5,000 units used 72,100 pounds, which were purchased at $11.27 per pound, what is the direct materials (a) price variance, (b) quantity variance, and (c) cost variance? Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

a. Direct materials price variance $
b. Direct materials quantity variance $
c. Direct materials cost variance $
Answer #1

a. Direct material price variance = (standard price - actual price)* actual quantity

= (11.5-11.27)*72100 = $16583 favorable

b. Direct material quantity variance = (5000*14 - 72100)*11.5 = 24150 unfavorable

c. Direct material cost variance = 5000*14*11.5 - 72100*11.27 = 7567 unfavorable